Kelso Corporation just paid a dividend of D_0 = $1.50 per share, and that dividend is expected to grow at a constant rate of 6.50% per year in the future. The company’s beta is 1.70, the required return on the market is 10.50%, and the risk-free rate is 4.50%. What is the company’s current stock price?
Answer #1

Required rate = Risk free rate + Beta *(Market return - risk Free rate) =4.50%+1.70*(10.50%-4.50%) = 14.70%

Current stock Price = D0*(1+Growth)/(Required Rate - growth) =1.50*(1+6.50%)/(10.50%-4.50%) = 26.63

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well